Workflow Team Meeting - Dec 11 4PM CERN time

Vidyo Link

Attending

Personel

Dec 11 -> Dec 17 Sara
  • EU - Do we have shifters over the Holidays?
  • Holiday vacation plans
    • Julian will be in Colombia 25th Nov - 25th Dec - Working plans on being online. and will be on holiday again Jan 1st-7th
    • Luis will be in Colombia Dec 20-through New Year will get us exact dates soon
      • Traveling on Dec 18, back Jan 5
    • Jen will be in MN Dec 26-Jan 2 and will have limited internet access
    • Ian & Sean will be around but nothing prolonged
    • Alan Dec 13-New Years Brazil - if you need him you have to call Alan's Grandmother, he will post number in Twiki wink
    • Seangchan will take a couple days off, but not telling us when wink
    • Dima will be in FL next week, but will be around over the Holidays
    • John (site support) Traveling on Sun 21, in Ecuador till end of Jan. Off until Jan 5, working remotely Jan 5-end of Jan
    • Jorge - Traveling Dec 19 - Jan 5 back on 6th not planning on working remotely - transfers
    • Andrew L(a) will be around (officially back on 5th Jan)

News

  • Ajit will help a bit watching things during the holidays, but he needs still to catch up.
  • Next 2 wks will be virtual due to Christmas and New Years! Relax and enjoy the Christmas Production :P
  • Because we will not be having meetings making sure we keep the chatter up on issues we are seeing in elog and actually reading e-logs mail will be extra important!

EU Operators Meeting Notes

  • we need more EU operators, and the Belgiums are having issues getting people setup
    • Sara - continuing, but working on her PhD but will not be around probably around first 6 months but not 2nd
    • Jasper - will not be available
    • Xavier - really busy
    • No new students coming in for next year so they will no longer be able to have a full person credit, we need to find a new institute to give us time.
  • None showed up to the meeting!
  • Division of responsibilities
    • Check wmagent/components up
    • Check workflows with high error %
    • Check sites with low running/pending %
  • Credits and stuff. - new operators. Julian will work with Christoph to get operators for other institutes
  • Campaigns going on: RunIIFall14GS, Phys14DR, Sprin14miniaod
  • Julian can be back watching things EU time after t he 27th
  • Nobody watching system EU time next week, Sara is traveling but we can ask directly what days she's available.

Site support

  • FNAL out of drain, please inform if any issues appear.
  • Sites to test: T2_RU_INR, T2_EE_Estonia
  • Sites temporarily in WR: T2_US_Nebraska, T2_PT_NCG_Lisbon, T2_IN_TIFR

Agent Issues

  • Proxy issue on Submit2 - where do we stand? Is Alan really the only one who can keep the proxies up?
    • Dave is planning on putting his proxy in place, log collect jobs will fail, but they are failing anyway. What is plan B for the Christmas Production if the Proxies on the SL6 machines can't be updated until we can get Alan's attention?
    • We are asking SeangChan and Dave to update the proxy and get it /keep it working
      • Seangchan will reply to the email thread - if we need to run any command or line.
      • We need a production - central proxy instead of a personal proxy.

Redeployment plan

* Upgrade condor on submit1 - only job. Jen will tell Krista.

Workflows

  1. TP2023HSCALGS and TP2023HGCALGS (Gen-sim)
  2. Fall14DR and Fall14DR73 (Digi Reco)
  3. Phys14DR for HLT and Summer12DR for Gen validation (Digi Reco)
  4. Phys14DR for HLT (Digi reco) RunIIFall14GS (Gen sim)
  5. Data Reco 2013A - Summer12DR and Phys14DR for Physics.
  6. Summer11LegDR - Summer12DR (Digi Reco) Summer12 - Summer11 (Gen Sim)

ReDigi

Waiting to be returned to requestor: https://cms-logbook.cern.ch/elog/Workflow+processing/18059 alahiff_BTV-Phys14DR-00006_00018_v0__141129_234223_67 pdmvserv_BTV-Phys14DR-00026_00041_v0__141112_153935_4359 pdmvserv_HIG-Phys14DR-00006_00031_v0__141111_144303_536 pdmvserv_BTV-Phys14DR-00010_00018_v0__141110_011817_633 pdmvserv_B2G-Spring14miniaod-00087_00075_v0__141030_150442_6231 pdmvserv_MUO-Spring14miniaod-00019_00084_v0__141030_155616_5795 pdmvserv_EXO-Spring14miniaod-00215_00077_v0__141030_153350_2462 pdmvserv_SUS-Spring14miniaod-00058_00072_v0__141030_100336_4528 pdmvserv_EXO-Spring14miniaod-00192_00068_v0__141030_093231_949 pdmvserv_EXO-Spring14miniaod-00163_00067_v0__141030_093040_4447 jen_a_BTV-Phys14DR-00003_00041_v0__141126_052704_5345

Miniaod with 100% error needs return to requestor:https://cms-logbook.cern.ch/elog/Workflow+processing/18049 pdmvserv_HIG-Spring14miniaod-00153_00098_v0__141120_134507_451

Bunny/Test WF's with 100% failure alahiff_BTV-Phys14DR-00011_00033_v0_castor_141211_105824_3160 alahiff_BTV-Phys14DR-00012_00033_v0_castor_141211_105833_422 alahiff_BTV-Phys14DR-00021_00033_v0_castor_141211_105851_2809 alahiff_BTV-Phys14DR-00023_00033_v0_castor_141211_105734_2400 alahiff_BTV-Phys14DR-00023_00033_v0_castor_141211_105900_2971 alahiff_EXO-Phys14DR-00078_00053_v0__141211_105743_7399 alahiff_EXO-Phys14DR-00078_00053_v0__141211_105752_6212 alahiff_EXO-Phys14DR-00109_00055_v0__141211_105801_3912 alahiff_EXO-Phys14DR-00109_00055_v0__141211_105815_5195 alahiff_BTV-Phys14DR-00016_00033_v0_castor_141211_105841_9406

Workflows not 100% dispite no errors except Log Collect : https://cms-logbook.cern.ch/elog/Workflow+processing/18058 alahiff_TSG-Phys14DR-00022_00032_v0__141129_233219_5496

WF's witn "None" as acquisition era :https://cms-logbook.cern.ch/elog/Workflow+processing/18105 pdmvserv_TOP-Summer12DR53X-00276_00354_v0__141202_101038_339 pdmvserv_TOP-Summer12DR53X-00275_00355_v0__141203_105458_7439 pdmvserv_HIG-Summer12DR53X-02171_00353_v0__141202_003657_5357

  • Dima is getting some answers.

miniaod's

  • still waiting for answer for the 100% failed wfs.

Rereco

Store Results

MonteCarlo

* Like 300 requests injected (RunIIFall14GS)! plenty of work for a while.

SL6 testing/SL5 Decomissioning

RelVal Andrew

  • made a pull request needs to talk to Seangchan and Alan, they are holding the merging until validation is done and then will do the pull after Tues.
  • We need to configure WMAgent to inject them into PhEDEx instance

AOB

-- JulianBadillo - 2014-12-17
Edit | Attach | Watch | Print version | History: r9 < r8 < r7 < r6 < r5 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r9 - 2014-12-18 - AndrewLahiff
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    CMSPublic All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2020 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
or Ideas, requests, problems regarding TWiki? use Discourse or Send feedback